# Separable Mixture of Low-Rank Adaptation for Continual Visual Instruction Tuning

### v1

@@ Abstract
Visual instruction tuning (VIT) enables multimodal large language models (MLLMs) to effectively handle a wide range of vision tasks by framing them as language-based instructions. Building on this, continual visual instruction tuning (CVIT) extends the capability of MLLMs to incrementally learn new tasks, accommodating evolving functionalities. While prior work has advanced CVIT through the development of new benchmarks and approaches to mitigate catastrophic forgetting, these efforts largely follow traditional continual learning paradigms, neglecting the unique challenges specific to CVIT. We identify a dual form of catastrophic forgetting in CVIT, where MLLMs not only forget previously learned visual understanding but also experience a decline in instruction following abilities as they acquire new tasks. To address this, we introduce the Separable Mixture of Low-Rank Adaptation (SMoLoRA) framework, which employs separable routing through two distinct modules—one for visual understanding and another for instruction following. This dual-routing design enables specialized adaptation in both domains, preventing forgetting while improving performance. Furthermore, we propose a novel CVIT benchmark that goes beyond existing benchmarks by additionally evaluating a model’s ability to generalize to unseen tasks and handle diverse instructions across various tasks. Extensive experiments demonstrate that SMoLoRA outperforms existing methods in mitigating dual forgetting, improving generalization to unseen tasks, and ensuring robustness in following diverse instructions.
@@ 1 Introduction
Visual instruction tuning for MLLMs is traditionally performed within a static multitask learning framework [5, 23], where a unified image-instruction-output data format is used. However, in real-world applications, MLLMs must be capable of learning new vision-language tasks to accommodate evolving requirements. This has sparked growing interest in continual visual instruction tuning (CVIT), which allows MLLMs to incrementally learn new tasks, reducing the significant costs associated with retraining models from scratch. Despite its potential, existing studies on continual learning have demonstrated that sequential fine-tuning can lead to catastrophic forgetting, where models forget previously learned knowledge when updated with new tasks. While recent work on CVIT [12, 3] has made strides in developing new benchmarks and strategies to mitigate forgetting, these efforts largely follow traditional continual learning paradigms and overlook the distinct challenges posed by CVIT’s integration of vision and language tasks.
In this work, we identify a dual form of catastrophic forgetting in CVIT. First, the model often loses its previously acquired visual understanding when updated with new tasks (as shown in Fig. 1 (a)). Second, its ability to follow instructions deteriorates as more tasks are learned over time (Fig. 1 (b)). Mixture-of-LoRA (MoLoRA) [24], originally developed for multitask learning, has been applied to CVIT [3]. MoLoRA uses multiple LoRA blocks coupled with a router network to capture task-specific knowledge, aiming to reduce forgetting. However, we find that this approach is less effective. The challenge in extending MoLoRA from multitask learning to continual learning lies in the complexity of dynamic routing. In multitask learning, tasks typically share common components, enabling efficient routing based on task similarity. In contrast, continual learning involves a constantly evolving stream of tasks with diverse, and sometimes conflicting, data distributions. This variability complicates the routing process, as the model must not only accommodate new tasks but also preserve previously learned knowledge from being overwritten or degraded. Without a proper routing strategy, new tasks can interfere with the retention of older knowledge, exacerbating the dual forgetting problem in CVIT.
Figure 2: Differences among LoRA-based methods for CVIT. Unlike SeqLoRA and MoLoRA, our proposed SMoLoRA utilizes separable routing and an adaptive fusion approach to enhance LoRA routing, aiming to mitigate dual catastrophic forgetting.
@@
We propose Separable Mixture-of-LoRA (SMoLoRA), a new approach specifically designed to tackle the dual forgetting problem in CVIT. SMoLoRA introduces separable routing, implemented through two distinct modules: a visual understanding module and an instruction following module. As illustrated in Fig. 2, unlike MoLoRA, which relies solely on input tokens to infer task-specific information, SMoLoRA’s routing strategy leverages both visual understanding and instruction following capabilities to more accurately assess the nature of the input. This dual-domain approach enables specialized adaptation in each domain, allowing the model to preserve and refine both its visual and instruction following abilities without interference. By dynamically selecting LoRA blocks tailored to the specific needs of each task, SMoLoRA effectively mitigates the dual forgetting problem. Our contributions are summarized as follows:
@@
  • We uncover the issue of dual catastrophic forgetting in CVIT and introduce the SMoLoRA framework as a solution. By leveraging separable routing, SMoLoRA dynamically selects appropriate LoRA blocks tailored to the specific needs of each task, effectively preserving and refining both visual understanding and instruction following capabilities—thus mitigating dual forgetting.
@@
  • We propose a new CVIT benchmark that not only evaluates model accuracy and forgetting on standard tasks but also assesses the model’s ability to generalize to unseen tasks and handle diverse instructions across a wide range of tasks, offering a more comprehensive evaluation.
  • Through extensive experiments on our CVIT benchmark, we show that SMoLoRA outperforms existing methods in addressing dual forgetting, enhancing generalization to new tasks, and demonstrating robust instruction following across diverse scenarios.
@@ 3 Method
Figure 3: SMoLoRA is an efficient MoLoRA method designed for CVIT, which can replace the original LoRA module in existing MLLMs. SMoLoRA consists of three modules. The visual understanding and instruction following modules use a separable routing strategy to select the most suitable LoRA blocks for the current input, thus preventing interference between dissimilar tasks. The adaptive fusion module performs weighted fusion on the results after separable routing using trainable parameters. Through the collaboration of these modules, SMoLoRA effectively addresses the problem of catastrophic forgetting in CVIT.
@@ 3.1 Continual Visual Instruction Tuning
Given a pre-trained MLLM $f$, continual visual instruction tuning (CVIT) aims to incrementally fine-tune $f$ to learn a set of vision tasks $\mathrm{T}=\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$. At each time step $t$, the model is presented with a new task $\tau_{t}$ alongside a new dataset $\mathcal{D}_{t}$, to integrate this new knowledge into its existing capabilities. Typically, the dataset $\mathcal{D}_{t}$ follows the format $\{X^{ins},X^{vis},X^{ans}\}$, where $X^{ins},X^{vis},X^{ans}$ denote textual instruction input, visual input, and linguistic answer, respectively. However, when learning the new task $\tau_{t}$, previously seen data $\{\mathcal{D}_{i}\}_{i=1}^{t-1}$ is inaccessible, which introduces the challenge of dual catastrophic forgetting: 1) The model tends to forget previously learned visual understanding ability when updated with new data, and 2) Its ability to follow instructions gradually declines as additional tasks are learned. Next, we introduce an SMoLoRA approach to efficiently mitigate the dual catastrophic forgetting problem in CVIT.
@@ 3.2 Separable Mixture-of-LoRA
Extending MoLoRA from multitask learning to continual learning presents a complex dynamic routing challenge. In multitask learning, tasks often share common expert components, allowing for efficient routing based on task similarity. However, in continual learning, the model is exposed to an evolving stream of tasks with highly diverse and sometimes conflicting data distributions. This variability makes the task of selecting the right experts more nuanced and non-trivial. As new tasks emerge, the model must adapt its routing strategy to not only accommodate new data but also ensure that previously learned tasks are not negatively impacted. Otherwise, new tasks can overwrite or diminish the knowledge associated with previous ones, resulting in dual forgetting problem in the context of CVIT.
@@ 3.3 Separable Routing and Adaptive Fusion
$G^{vu}(z^{vu})=\operatorname{softmax}\left(\operatorname{top}_{k}\left(R^{vu}\text{Avg}(x_{l-1})\right)\right),$ (5)
@@
In contrast, the differences in instruction requirements are also a key factor contributing to the task discrepancies. The instruction following module uses the embedding of the current task instruction $f_{\sigma}(X^{ins})\in\mathbb{R}^{e\times 1}$ as the input $z^{if}$ for routing, which helps reduce interference among different task requirements during continual learning:
$G^{if}(z^{if})=\operatorname{softmax}\left(\operatorname{top}_{k}\left(R^{if}f_{\sigma}(X^{ins})\right)\right),$ (6)
@@
Adaptive fusion. Considering that the contributions of the two modules’ router strategies may not be identical during learning, we adopt an adaptive approach to assign weights to the outputs of these two modules. First, we compute the outputs of the two modules $x_{l}^{vu}$ and $x_{l}^{if}$ separately:
@@
$[\alpha,\beta]^{T}=\operatorname{softmax}\left(\operatorname{concat}\left(I^{vu}x_{l}^{vu},I^{if}x_{l}^{if}\right)\right),$ (8)
where $\operatorname{concat}(\cdot)$ represents concatenation along the first dimension. The final result after this adaptive fusion is given by:
@@
where $\circ$ represents the element-wise multiplication of matrices.
@@ 4.1 Benchmark Construction
Continual stability and transfer generalization. The learning and evaluation process of the CVIT benchmark is divided into two stages: upstream continual learning and downstream knowledge transfer. For the upstream tasks, we sequentially select ScienceQA [26], TextVQA [34], Flickr30k [29], ImageNet [6], GQA [16], and VQAv2 [10] as subjects for continual learning and stability evaluation. Regarding the downstream tasks, we evaluate the zero- and few-shot performance across four diverse datasets after completing all upstream tasks. VizWiz [11], TextCaps [33], and OCRVQA [27] are evaluated under zero-shot settings, whereas Places365 [44] is evaluated under both 5-shot and 10-shot settings.
Exploration of instruction diversity. To more thoroughly assess model adaptability and robustness, we incorporate both single-type and multi-type instructions into each task. This approach mirrors real-world scenarios and allows for a systematic analysis of the impact of instruction diversity on model performance. It also enables the evaluation of various CVIT methods’ adaptability under conditions involving both single and multiple instruction types.
4.2 Evaluation Metrics Design
@@
where $\mathcal{B}_{j}(o_{i}^{j})$ indicates whether the output $o_{i}^{j}$ of the $j$-th task conforms to the required instruction format for the current task, $n$ denotes the number of test samples. This is a binary classification evaluation.
@@ 5.1 Experiment Setup
Comparison methods. We compare our method against a diverse set of alternatives to highlight its superior performance. Multitask and Zero-Shot represent training all tasks jointly and no task-specific training, respectively. DirLoRA introduces a separate LoRA for each individual task, while SeqLoRA [14] sequentially fine-tunes all tasks using a single shared LoRA. We also include the enhanced LoRA variant, DoRA [25], for comparison. Additionally, we evaluate classic continual learning approaches such as EWC [19] and Replay [2]. Finally, we assess MoeLoRA [3], a token-wise MoLoRA method tailored for CVIT.
Implemented details. We use the pre-trained LLaVA-v1.5-7B [23] as the base model for our experiments. The hyperparameters for SMoLoRA are configured as follows: a learning rate of $1\times 10^{-4}$ with a cosine decay schedule, and a batch size of 64. Training is performed over a single epoch. The number of LoRA blocks is set to 4 for both visual understanding module and instruction following module, with each LoRA having a rank of 16, and $\operatorname{top}_{k}(\cdot)$ set to $\operatorname{top}_{1}(\cdot)$. Both SMoLoRA and the baseline methods are applied specifically to the feedforward network (FFN) layers of the LLM and the adapter.
@@
5.2 Main Results
Fig. 5 further illustrates the performance trends across the first four datasets throughout the continual instruction tuning process. It is clear that our method consistently sustains high performance during continual learning, while other approaches experience abrupt performance drops due to catastrophic forgetting.
5.3 Ablation Study
Effect of the components in SMoLoRA. We begin by examining the impact of each component in SMoLoRA on the single-type CVIT benchmark. The VU, IF, and AF modules represent the visual understanding, instruction following, and adaptive fusion components, respectively. Without any of these modules is equivalent to SeqLoRA. In contrast, when the VU or IF module is added, performance improves significantly, with the IF module yielding the most substantial gains, particularly in the MIF metric. This underscores the critical role of the instruction following in task learning within CVIT. When both the VU and IF modules are combined by averaging their outputs, performance improves further. The inclusion of the AF module provides an additional boost, emphasizing its effectiveness in enhancing the model’s overall capabilities.
@@ 5.3 Ablation Study
Effect of routing of LoRA blocks. To further explore how SMoLoRA mitigates the dual catastrophic forgetting issue, we randomly select a fixed number of examples from each dataset and visualize their LoRA blocks routing behaviors in Fig. 7. Longer color bars indicate a stronger preference for specific LoRA blocks. Within the VU module’s four LoRA blocks, we observe considerable variation in selections across different tasks. However, similar tasks, such as the captioning tasks of Flickr30k and TextCaps, show converging routing patterns. In the IF module, where routing is driven by task-specific instructions, tasks with differing instructional requirements demonstrate distinct preferences for different LoRA blocks. This visualization highlights the precise control our method exerts during the routing process, which effectively mitigates catastrophic forgetting problem by preventing improper LoRA block allocations across tasks.

### v2

# SMoLoRA: Exploring and Defying Dual Catastrophic Forgetting in Continual Visual Instruction Tuning

@@ Abstract
Visual instruction tuning (VIT) enables multimodal large language models (MLLMs) to effectively handle a wide range of vision tasks by framing them as language-based instructions. Building on this, continual visual instruction tuning (CVIT) extends the capability of MLLMs to incrementally learn new tasks, accommodating evolving functionalities. While prior work has advanced CVIT through the development of new benchmarks and approaches to mitigate catastrophic forgetting, these efforts largely follow traditional continual learning paradigms, neglecting the unique challenges specific to CVIT. We identify a dual form of catastrophic forgetting in CVIT, where MLLMs not only forget previously learned visual understanding but also experience a decline in instruction following abilities as they acquire new tasks. To address this, we introduce the Separable Mixture of Low-Rank Adaptation (SMoLoRA) framework, which employs separable routing through two distinct modules—one for visual understanding and another for instruction following. This dual-routing design enables specialized adaptation in both domains, preventing forgetting while improving performance. Furthermore, we propose a new CVIT benchmark that goes beyond existing benchmarks by additionally evaluating a model’s ability to generalize to unseen tasks and handle diverse instructions across various tasks. Extensive experiments demonstrate that SMoLoRA outperforms existing methods in mitigating dual forgetting, improving generalization to unseen tasks, and ensuring robustness in following diverse instructions. Code is available at https://github.com/Minato-Zackie/SMoLoRA.
@@ 1 Introduction
Visual instruction tuning for MLLMs is typically performed within a static multitask learning framework [5, 23], where a unified image-instruction-output data format is used. However, in real-world applications, MLLMs need to learn new vision-language tasks to accommodate evolving requirements. This practical necessity has sparked a growing interest in continual visual instruction tuning (CVIT), which allows MLLMs to incrementally learn new tasks, reducing the significant costs associated with retraining models from scratch. Despite its potential, existing studies [41, 12] on continual learning have demonstrated that sequential fine-tuning can lead to catastrophic forgetting, where models forget previously learned knowledge when updated with new tasks. While recent work [12, 3] has made strides in developing new benchmarks and approaches to mitigate forgetting, these efforts largely follow traditional continual learning paradigms and overlook the distinct challenges in CVIT.
In this work, we identify a dual form of catastrophic forgetting in CVIT. First, the model often loses its previously acquired visual understanding when updated with new tasks ( See Fig. 1 (a)). Second, its ability to follow instructions deteriorates as more tasks are learned over time (See Fig. 1 (b)). Both forms of forgetting lead to performance degradation during the CVIT process. Inspired by the mixture-of-expert approaches [31, 22] and LoRA-based tuning approaches [14, 25], we propose a new Separable Mixture-of-LoRA (SMoLoRA) approach, specifically designed to tackle the dual forgetting problem in CVIT.
@@
In SMoLoRA, we introduce a new separable routing mechanism by leveraging two specialized modules: a visual understanding module and an instruction following module. This dual-domain approach enables specialized adaptation in each domain, allowing the model to preserve and refine both its visual and instruction-following abilities without interference. By dynamically selecting LoRA blocks tailored to the specific needs of each task, SMoLoRA effectively mitigates the dual forgetting problem. Additionally, we introduce a new benchmark to comprehensively evaluate CVIT performance of various approaches, overcoming the limitations of existing benchmarks [12, 3], such as restricted task and instruction diversity, insufficient evaluation on unseen datasets, and the lack of comprehensive metrics.
Our contributions are summarized as follows:
@@
  • We uncover the issue of dual catastrophic forgetting in CVIT and introduce the SMoLoRA framework as a solution. Through a new separable routing mechanism, SMoLoRA dynamically selects task-specific LoRA blocks, optimizing both visual understanding and instruction-following capabilities, thus mitigating dual forgetting.
@@
  • We propose a new CVIT benchmark that not only evaluates task-specific performance but also emphasizes generalization to unseen tasks and the model’s ability to handle diverse instructions, providing a more holistic assessment.
  • Through extensive experiments on our CVIT benchmark, we demonstrate that existing approaches struggle with dual forgetting and generalization to new tasks. In contrast, SMoLoRA excels across a wide range of metrics and scenarios, delivering superior performance.
@@ 2 Related Work
Figure 2: SMoLoRA is an effective method designed for CVIT, which can replace the original LoRA module in existing MLLMs. SMoLoRA consists of three modules. The visual understanding and instruction following modules use a separable routing strategy to select the most suitable LoRA blocks for the current input, thus preventing interference between dissimilar tasks. The adaptive fusion module performs weighted fusion on the results after separable routing using trainable parameters. Through the collaboration of these modules, SMoLoRA effectively addresses the problem of dual catastrophic forgetting in CVIT.
@@ 3.1 Continual Visual Instruction Tuning
Given a pre-trained MLLM $f$, continual visual instruction tuning (CVIT) aims to incrementally fine-tune $f$ to learn a set of vision tasks $\mathrm{T}=\{\tau_{1},\tau_{2},\ldots,\tau_{n}\}$. At each time step $t$, the model is presented with a new task $\tau_{t}$ alongside a new dataset $\mathcal{D}_{t}$, to integrate this new knowledge into its existing capabilities. Typically, the dataset $\mathcal{D}_{t}$ follows the format $\{X^{ins},X^{vis},X^{ans}\}$, where $X^{ins},X^{vis},X^{ans}$ denote textual instruction input, visual input, and linguistic answer, respectively. However, when learning the new task $\tau_{t}$, previously seen data $\{\mathcal{D}_{i}\}_{i=1}^{t-1}$ is inaccessible, which introduces the challenge of dual catastrophic forgetting: 1) The model tends to forget previously learned visual understanding ability when updated with new data, and 2) Its ability to follow instructions gradually declines as additional tasks are learned. Next, we introduce an SMoLoRA approach to effectively mitigate the dual catastrophic forgetting problem in CVIT.
@@ 3.2 Separable Mixture-of-LoRA
Extending MoLoRA from multitask learning to continual learning presents a complex dynamic routing challenge. In multitask learning, tasks often share common expert components based on task similarity. However, in continual learning, the model is exposed to an evolving stream of tasks with highly diverse and sometimes conflicting data distributions. This variability makes the task of selecting the right experts more nuanced and non-trivial. As new tasks emerge, the model must adapt its routing strategy to not only accommodate new data but also ensure that previously learned tasks are not negatively impacted. Otherwise, new tasks can overwrite or diminish the knowledge associated with previous ones, resulting in dual forgetting problem in the context of CVIT.
@@ 3.3 Separable Routing and Adaptive Fusion
$G^{vu}(z^{vu})=\underset{\text{dim}=0}{\operatorname{softmax}}\left(\operatorname{top}_{k}\left(R^{vu}\text{Avg}(x_{l-1})\right)\right),$ (5)
@@
In contrast, the differences in instruction requirements also play a crucial role in contributing to task discrepancies. The instruction following module leverages the embedding of the current task instruction, extracted using Sentence-BERT [30], $f_{\sigma}(X^{ins})\in\mathbb{R}^{e\times 1}$, as the input $z^{if}$ for routing. This approach helps mitigate interference between different task requirements during continual learning:
$G^{if}(z^{if})=\underset{\text{dim}=0}{\operatorname{softmax}}\left(\operatorname{top}_{k}\left(R^{if}f_{\sigma}(X^{ins})\right)\right),$ (6)
@@
Adaptive fusion. Considering that the contributions of the two modules’ router strategies may not be identical during learning, we adopt an adaptive approach to assign weights to the outputs of these two modules. First, we compute the outputs of the two modules $x_{l}^{vu}$ and $x_{l}^{if}\in\mathbb{R}^{k\times s}$ separately:
@@
$[\alpha,\beta]^{T}=\underset{\text{dim}=-1}{\operatorname{softmax}}\left(\operatorname{concat}\left(I^{vu}x_{l}^{vu},I^{if}x_{l}^{if}\right)\right),$ (8)
where ${\alpha,\beta\in\mathbb{R}^{1\times s}}$, $\operatorname{concat}(\cdot)$ represents concatenation along the first dimension. The final result is given by:
@@
where $\circ$ represents the element-wise multiplication applied independently to each row of the matrices.
@@ 4.1 Benchmark Construction
Continual stability and transfer generalization. The CVIT benchmark is used in a two-stage manner: upstream continual learning and downstream knowledge transfer. For the upstream tasks, we sequentially select ScienceQA [26], TextVQA [35], Flickr30k [29], ImageNet [6], GQA [16], and VQAv2 [10] as subjects for continual learning and stability evaluation. Regarding the downstream tasks, we evaluate the zero- and few-shot performance across four diverse datasets after completing all upstream tasks. VizWiz [11], TextCaps [34], and OCRVQA [27] are evaluated under zero-shot settings, whereas Places365 [46] is evaluated under both 5-shot and 10-shot per-class settings.
Exploration of instruction diversity. To more thoroughly assess model adaptability and robustness, we incorporate both single-type and multi-type instructions into each task, which mirrors real-world scenarios and allows for a systematic analysis of the impact of instruction diversity on model performance. It also enables evaluating various CVIT methods’ adaptability under conditions involving both instruction types. More details about CVIT benchmark are provided in Section D of the supplementary material.
4.2 Metrics Designing
@@
where $\mathcal{B}()$ is a binary decision function that outputs 1 if the result $o_{i}^{j}$ matches the current task’s instruction format, and 0 otherwise. $n$ denotes the number of test samples.
@@ 5.1 Experiment Setup
Comparison methods. We compare our method against a diverse set of alternatives to highlight its superior performance. Multitask and Zero-Shot represent training all tasks jointly and no task-specific training, respectively. DirLoRA introduces a separate LoRA for each individual task, while SeqLoRA [14] sequentially fine-tunes all tasks using a single shared LoRA. We also include the enhanced LoRA variants, DoRA [25] and C-LoRA [36], for comparison. Additionally, we evaluate classic continual learning approaches such as EWC [19] and Replay [2]. Finally, we assess a series of methods specifically designed for CVIT, including EWC+TIR [12], Eproj [12], and a token-wise MoLoRA method, MoeLoRA [3].
Implemented details. We use the pre-trained first-stage LLaVA-v1.5-7B [23] as our base model. The hyperparameters for SMoLoRA are configured as follows: a learning rate of $1\times 10^{-4}$ with a cosine decay schedule, and a batch size of 64. Training is performed over a single epoch. The number of LoRA blocks is set to 4 for both visual understanding module and instruction following module, with each LoRA having a rank of 16, and $\operatorname{top}_{k}(\cdot)$ set to $\operatorname{top}_{1}(\cdot)$. Both SMoLoRA and the baseline methods are applied specifically to the feedforward network (FFN) layers of the LLM and the adapter. We also conduct relevant experiments on MiniGPT-4 [47], as detailed in supplementary material C.
@@
5.2 Main Results
Fig. 4 further illustrates the performance trends across the first four datasets throughout the continual instruction tuning process. It is clear that our method consistently sustains high performance during continual learning, while other approaches experience abrupt performance drops due to catastrophic forgetting. See supplementary material B for performance variation details.
5.3 Ablation Study
Effect of the components in SMoLoRA. We begin by examining the impact of each component in SMoLoRA on the single-type CVIT benchmark. The VU, IF, and AF modules represent the visual understanding, instruction following, and adaptive fusion components, respectively. Without any of these modules is equivalent to SeqLoRA. In contrast, when the VU or IF module is added, performance improves significantly, with the IF module yielding the most substantial gains, particularly in the MIF metric. This underscores the critical role of the instruction following in task learning within CVIT. When both the VU and IF modules are combined by averaging their outputs, performance improves further. The inclusion of the AF module provides an additional boost, emphasizing its effectiveness in enhancing the model’s overall capabilities. We further show how the AF module balances the proportional relationship between the VU and IF modules in supplementary material A.
@@ 5.3 Ablation Study
Effect of routing of LoRA blocks. To further explore how SMoLoRA mitigates the dual catastrophic forgetting issue, we randomly select a fixed number of examples from each dataset and visualize their LoRA blocks routing behaviors in Fig. 6. Longer color bars indicate a stronger preference for specific LoRA blocks. Within the VU module’s four LoRA blocks, we observe considerable variation in selections across different tasks. However, similar tasks, such as the captioning tasks of Flickr30k and TextCaps, show converging routing patterns. In the IF module, where routing is driven by task-specific instructions, tasks with differing instructional requirements demonstrate distinct preferences for different LoRA blocks. This visualization highlights the precise control our method exerts during the routing process, which effectively mitigates forgetting by preventing improper LoRA block allocations across tasks.
@@
Figure 7: Case study. Red text indicates inappropriate outputs, while blue text represents appropriate outputs.
5.4 Case Study
Finally, we present a case study in Fig. 7 to illustrate how SMoLoRA effectively addresses dual forgetting issues through the integration of two specialized modules: the Visual Understanding (VU) module, designed to mitigate visual understanding forgetting, and the Instruction Following (IF) module, aimed at tackling instruction following forgetting. When only the VU module is applied, the model comprehends visuals but fails to generate coherent sentences, producing fragmented words instead. Conversely, with only the IF module, the model adheres to the required text format but misinterprets visual targets. Integrating both modules achieves robust performance for both visual understanding and instruction adherence. Additional case studies can be found in Section E of the supplementary material.